OpenStudy (anonymous):

Judith has two closed rectangular prism recycling bins. The dimensions of the second recycle bin are one-fifth of the dimensions of the first recycle bin. Which statement is true about the surface area of the second recycle bin?

OpenStudy (lgbasallote):

volume of first prism = length x width x height so if you take the 1/5 of each volume of second prism = 1/5 length x 1/5 width x 1/5 height volume = (1/5 x 1/5 x 1/5)(length x width x height) find that 1/5 x 1/5 x 1/5 that's your answer